Determine whether each of the following points is inside, outside, or on the circle with equation x^2+y^2=5.
A) (-1,2)
B) (3/2,2)
C) (-1,3/2)
1
Expert's answer
2020-08-24T18:33:31-0400

ANSWER: point A) is located on the circle, point B)is located outside the circle, point C)is located inside the circle

EXPLANATION

A) "(-1,2)\quad : { (-1) }^{ 2 }{ +2 }^{ 2 }=1+4=5"

B) "(3/2,2):\ \ { (3/2) }^{ 2 }\quad { +2 }^{ 2 }=9/4+4=25/4>5"

C)"(-1,3\/2): { (-1) }^{ 2 }+{ (3\/2) }^{ 2 }=1+9\/4=13\/4<5"